The cooked up spread. You have been doing research on Ford Motor Company (F). You note...

  1. The cooked up spread. You have been doing research on Ford Motor Company (F). You note that Ford is currently trading at $9.35. After doing some analysis you feel that a fair price range over the next year for Ford is between $15 and $17. You however, are unsure if you wish to invest in Ford over a long horizon and feel that an options trade may be most appropriate. You see that there are listed options with a June 26th expiration and that these would work for your purposes. You decide that you want to long one call with a $10 strike, you also decide that you don’t want to spend much money entering the trade, so you short 10 calls with a strike of $15 as this aligns with your price range on F. What is the dollar outlay on this trade, max profit, and max loss (3 pts, one point per answer)? In your opinion is the potential profit with the risk (one point)? Now, what if instead you also buy 9 calls with a strike of $17 which was on the high side of your price estimate. What would be the max loss, max profit, and the two break even points if this is the trade you set up (4 points, one point per)?
  2. 5.00

    4.40
    8.00 1.59
    10.00 0.52
    12.00 0.15
    15.00 0.05
    17.00 0.03
    20.00 0.02
    22.00 0.02
Answer #1
1.Cash Flow for Buying (Long)one Call at strike $10 -$0.52
2.Cash Flow for Selling (Short) Ten Call at strike$15 $0.50 (0.05*10)
Net Cash Flow -$0.02
Assume Price at expiration =S
Payoff for One Long Call at Strike =$10
Payoff:Max.((S-10),0)
Payoff for One Short Call at Strike =$15
Payoff:Min.((15-S),0)
A B C=A+B-0.02
Payoff Payoff Net
Price at Expiration 1 Long Call $10 10 Short Call $15 Profit
$5 $0 $0 ($0.02)
$6 $0 $0 ($0.02)
$7 $0 $0 ($0.02)
$8 $0 $0 ($0.02)
$9 $0 $0 ($0.02)
$10 $0 $0 ($0.02)
$10.02 $0.02 $0 ($0.00)
$11 $1 $0 $0.98
$12 $2 $0 $1.98
$13 $3 $0 $2.98
$14 $4 $0 $3.98
$15 $5 $0 $4.98
$16 $6 ($10) ($4.02)
$17 $7 ($20) ($13.02)
$18 $8 ($30) ($22.02)
$19 $9 ($40) ($31.02)
$20 $10 ($50) ($40.02)
$21 $11 ($60) ($49.02)
$22 $12 ($70) ($58.02)
$23 $13 ($80) ($67.02)
$24 $14 ($90) ($76.02)
$25 $15 ($100) ($85.02)
Dollar Outlay $0.02
Maximum Profit $4.98
Maximum Loss Infinite
Break Even Point $10.02
ADD:
BUY (Long)9 calls with Strike $17
Payoff :Max.((S-17),0) for One Long Call at Strike =$17
Cash Flow =9*0.03 ($0.27)
Net Cash Flow =-0.02-0.27 ($0.29)
A B C D=A+B+C-0.29
Payoff Payoff Payoff Net
Price at Expiration 1 Long Call $10 10 Short Call $15 9 Long Call $17 Profit
$5 $0 $0 $0 ($0.29)
$6 $0 $0 $0 ($0.29)
$7 $0 $0 $0 ($0.29)
$8 $0 $0 $0 ($0.29)
$9 $0 $0 $0 ($0.29)
$10 $0 $0 $0 ($0.29)
$10.29 $0.29 $0 $0 ($0.00)
$11 $1 $0 $0 $0.71
$12 $2 $0 $0 $1.71
$13 $3 $0 $0 $2.71
$14 $4 $0 $0 $3.71
$15 $5 $0 $0 $4.71
$16 $6 ($10) $0 ($4.29)
$17 $7 ($20) $0 ($13.29)
$18 $8 ($30) $9 ($13.29)
$19 $9 ($40) $18 ($13.29)
$20 $10 ($50) $27 ($13.29)
$21 $11 ($60) $36 ($13.29)
$22 $12 ($70) $45 ($13.29)
$23 $13 ($80) $54 ($13.29)
$24 $14 ($90) $63 ($13.29)
$25 $15 ($100) $72 ($13.29)
Dollar Outlay $0.29
Maximum Profit $4.71
Maximum Loss ($13.29)
Break Even point $10.29